A body of mass 300g is accelerated uniformly from rest across an horizontal surface and covers a distance of 50m, in 20s . calculate

a) The magnitude of the acceleration of the body across a surface

b) The magnitude of the minimum force which produces the acceleration obtained in (a) above and

c) The moment of the body immediately after covering the 50m


1
Expert's answer
2021-03-16T09:05:05-0400

a)


"s=at^2/2\to a=2s/t^2=2\cdot50/20^2=0.25\ (m/s^2)"


b)


"F=ma=0.3\cdot0.25=0.075\ (N)"


c)


"p=mv=m\cdot at=0.3\cdot0.25\cdot20=1.5\ (kg\cdot m/s)"